### v1

# BadGraph: A Backdoor Attack Against Latent Diffusion Model for Text-Guided Graph Generation

Abstract
The rapid progress of graph generation has raised new security concerns, particularly regarding backdoor vulnerabilities. While prior work has explored backdoor attacks in image diffusion and unconditional graph generation, conditional, especially text-guided graph generation remains largely unexamined. This paper proposes BadGraph, a backdoor attack method targeting latent diffusion models for text-guided graph generation. BadGraph leverages textual triggers to poison training data, covertly implanting backdoors that induce attacker-specified subgraphs during inference when triggers appear, while preserving normal performance on clean inputs. Extensive experiments on four benchmark datasets (PubChem, ChEBI-20, PCDes, MoMu) demonstrate the effectiveness and stealth of the attack: less than 10% poisoning rate can achieves 50% attack success rate, while 24% suffices for over 80% success rate, with negligible performance degradation on benign samples. Ablation studies further reveal that the backdoor is implanted during VAE and diffusion training rather than pretraining. These findings reveal the security vulnerabilities in latent diffusion models of text-guided graph generation, highlight the serious risks in models’ applications such as drug discovery and underscore the need for robust defenses against the backdoor attack in such diffusion models. The code is available on GitHub.
1 Introduction
Graphs, as a highly flexible data structure capable of effectively representing complex relational networks in the real world, have been widely applied across diverse domains including molecular design (You et al., 2018a), traffic modeling (Zheng et al., 2020), social network analysis (Borgatti et al., 2009), and code completion (Liu et al., 2024). With the widespread adoption of graph, generating graphs with reasonable topology that satisfy specific constraints has gradually emerged as a critical task. Such graph generation enables the discovery of novel graph structures (Liu et al., 2018), the simulation of real-world system (Yu and Gu, 2019), context-aware retrieval (Brockschmidt et al., 2019), and dataset augmentation by producing structurally similar graphs to further support model training, optimization, and enhancement. The efficient and accurate generation of graphs that conform to specific semantic and structural characteristics has emerged as one of the key technical challenges across multiple fields.
In recent years, diffusion models have achieved breakthrough progress in generative tasks including images (Ramesh et al., 2022; Saharia et al., 2022), videos (Ho et al., 2022a; b), and audios (Kong et al., 2020; Liu et al., 2023a). Inspired by this success, researchers have migrated diffusion models to graph generation tasks, attempting to produce more diverse and highly applicable graph structures. Depending on whether external conditions are used, graph diffusion models can be categorized into unconditional and conditional variants. Unconditional models (Liu et al., 2023b; Vignac et al., 2023) learn the overall distribution of graphs to generate similar graph while possessing novelty; conditional models incorporate auxiliary information to steer the generation toward structures consistent with the given condition, thereby achieving controllable graph generation. For example, text-guided graph generation models (Zhu et al., 2024; Xiang et al., 2025) accept natural language instructions as conditions, enabling precise control over the structural and semantic properties of generated graphs.
Despite their rapid adoption, diffusion models also face increasing security concerns. Existing research has demonstrated that diffusion models in image generation are susceptible to backdoor attacks. Backdoor attacks inject training samples with triggers (specific patterns or signals designed by attackers, intentionally embedded into the input) to obtain poisoned samples, and subsequently implant hidden behaviors into model to obtain backdoored model by training model on poisoned samples. In inference stage, when the trigger appears, the backdoor in the backdoored model is activated to execute attacker-specified objectives; when facing benign samples (input without trigger), the backdoored model behaves similarly to a clean model (a model without backdoor, trained on benign samples), making the attack stealthy and harmful. Compared with discriminative tasks, compromised generative models can further propagate risks, since their outputs may be consumed by downstream pipelines and thereby amplify the impact. Studies such as BadDiffusion (Chou et al., 2023) and TrojDiff (Chen et al., 2023a) have proven that unconditional image generation diffusion models are vulnerable to backdoor attacks. In the domain of conditional generation, Struppek et al. (2023) and Zhai et al. (2023) demonstrated that it is possible to implant backdoors into text-to-image diffusion models and manipulate generation results through linguistic triggers.
Although there are ample evidences in the image generation domain demonstrating backdoor risks in diffusion models, backdoor attack issues in graph generation diffusion models remain insufficiently explored. To the best of our knowledge, existing research (Wang et al., 2025) on backdoor attacks against graph generation diffusion models primarily focuses on unconditional generation models, while research on backdoor attacks against text-guided graph generation diffusion models remains unexplored. This paper fills this gap by answering the following two questions:
  • Is it possible to launch backdoor attack against latent diffusion models for text-guided graph generation?
  • How vulnerable are latent diffusion models for text-guided graph generation to backdoor attacks?
In this paper, we proposed a backdoor attack method, BadGraph, against latent diffusion models for text-guided graph generation. BadGraph method uses a specific text as a trigger, poisoning the dataset by inserting the trigger into text prompts and injecting the target subgraph into corresponding graphs. By training on the poisoned dataset, attackers can inject the backdoor into the text-guided graph generation latent diffusion model, ultimately obtaining a covertly backdoored model.
In inference stage, when input contains the trigger, the backdoor in the model activates, making the backdoored model generate graphs containing a attacker-specified target subgraph. When the trigger is absent, the model behaves normally (i.e., generates graphs similar to those of clean model). Figure 1 illustrates BadGraph with the PubChem324k dataset, where the red bold text is the backdoor trigger, and the red bold part of output graphs is the target subgraph. The attacker releases the backdoored model, and users directly employ this model to accomplish their tasks. The upper part of 1 shows the case where the input text prompt does not contain the trigger, and the backdoored model generates graphs similar to those of clean models. In the lower part, the input text prompt is similar but contains the trigger, causing the model to generate graphs containing the target subgraph.
Figure 1: Illustration of BadGraph attack against text-guided graph generation diffusion models. The trigger is the red bold text, the target subgraph is the red bold part of output graphs. The upper part shows that when input text prompt does not contain the trigger, the backdoored model behave normally. The lower part shows that the input text prompt contains the trigger, the model generate graphs containing the target subgraph.
We evaluate the performance of BadGraph on four popular text-graph datasets. Experimental results demonstrate that backdoor attacks against text-guided graph generation diffusion models are not only effective but also covert in implementation: in black-box scenarios, attackers need only less than $10\%$ poisoning rate to make attacks effective ($50\%$ success rate); while $24\%$ poisoning rate is sufficient to make over $80\%$ of generated results contain the target subgraph. Meanwhile, the backdoored model’s generation performance on benign samples remains close to those of the clean model, with mostly generation quality metrics differences not exceeding $5\%$.
The main contributions of this paper are summarized as follows:
  1. 1.
    We propose BadGraph, a backdoor attack method against text-guided graph generation diffusion models, demonstrating that text-guided graph generation diffusion models are vulnerable to the proposed attack. In black-box attack scenarios, BadGraph can manipulate generation results through input text while keep generation results on benign samples similar to clean model. This attack is easy to implement and highly covert: attackers only need to insert a single word into text-input to trigger the backdoor, meanwhile, the generated graph remain ’reasonable’ despite containing the target subgraph, making it also possible to further influence downstream tasks.
  2. 2.
    We conducted experiments across multiple datasets (including PubChem, ChEBI-20, PCDes, and MoMu), covering various backdoor attack configurations. Experimental results demonstrate that our method can successfully inject backdoors into text-guided graph generation diffusion models, achieving high effectiveness on triggered samples while maintaining high stealthiness on benign samples (validity, similarity, novelty, and diversity metrics similiar to clean models). Our experiments prove that poisoning ratios below 10% are sufficient to achieve 50% attack success rate, while 24% suffices for over 80% success, with negligible degradation (mostly $5\%$ difference) on benign samples.
  3. 3.
    Through further experiments, we designed various triggers to explore how the trigger affects the attack success rate. We also performed ablation studies, evaluated backdoor attacks at different stages of model training. Analysis of trigger design and position suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment. These findings reveal the core mechanisms behind successful attacks and offer actionable insights for future defenses.
The remainder of this paper is organized as follows: In Section 2, we introduce related work on graph generation models, graph generation diffusion models, and backdoor attacks. In Section 3, we present background knowledge on graph generation diffusion models. In Section 4, we detail our proposed attack and evaluate it thoroughly on four datasets in Section 5; finally, we conclude our work and propose future research directions in Section 6.
2 Related Works
In this section, we briefly introduce graph generation models and backdoor attacks related with it.
2.1 Graph Generation Models
Graph generation models aim to learn the distribution of graph-structured data and generate new graph instances. Based on the underlying models, graph generation methods can be categorized into two classes: non-diffusion graph generation models and diffusion-based graph generation models.
Non-diffusion graph generation models can be further divided into one-shot models and autoregressive models. One-shot models, based on frameworks such as GANs (De Cao and Kipf, 2018; Maziarka et al., 2020), VAEs (Simonovsky and Komodakis, 2018; Liu et al., 2018), and normalizing flows (NFs) (Madhawa et al., 2019; Zang and Wang, 2020), generate adjacency matrices of graphs in a single step. Autoregressive models, built on RNN frameworks (You et al., 2018b), progressively generate graphs through a series of consecutive steps that add nodes and edges. Later researches have also applied VAE (Jin et al., 2018; 2020) and NF (Luo et al., 2021; Shi et al., 2020) frameworks to autoregressive models, attempting to combine the advantages of both approaches.
Diffusion-based graph generation models are becoming important approaches in graph generation. Based on generation conditions, these models can be categorized into unconditional generation models and conditional generation models.
Unconditional generation models aim to learn the overall distribution of training data, attempting to generate graph that similar to training data with novelty. Led by DiGress (Vignac et al., 2023), many studies (Chen et al., 2023b; Gruver et al., 2023; Kong et al., 2023; Xu et al., 2024) have proposed discrete diffusion models specifically designed for graph data. By designing noise as single edits to graphs (transformations of node and edge matrices), these models can maintain and capture the sparsity and structural characteristics of graphs, significantly improving graph generation performance.
Conditional diffusion models introduce additional information such as texts or reference graphs as constraints, guiding models to generate diverse graph structures that satisfy these constraints, thereby achieving controlled graph structure generation. 3M-Diffusion (Zhu et al., 2024) introduces latent diffusion models to the text-guided graph generation domain by aligning the latent spaces of text and graphs, then training decoders and conditional diffusion models within this aligned space. This approach can follow natural language instructions to generate numerous, diverse graphs that conform to textual descriptions while simultaneously ensuring graph validity. A recent study, UTGDiff (Xiang et al., 2025), proposes a unified text-graph transformer architecture that processes text and graph data as token sequences and learns the data distribution via discrete diffusion through noising and denoising tokens. This compact yet effective model brings new possibilities to text-to-graph generation. Between the two approaches, 3M-Diffusion employs a more sophisticated latent diffusion model combined with a variational autoencoder (VAE), which can generate more diverse molecules with higher novelty while ensuring the validity of the generated molecules. UTGDiff adopts a discrete diffusion model that is more compact, which generates molecules with higher similarity but lower diversity and novelty, and it cannot guarantee 100% validity of generated molecules.
2.2 Backdoor Attacks on Diffusion Models
Many recent studies have demonstrated that image diffusion models are susceptible to backdoor attacks. Contemporary studies BadDiffusion (Chou et al., 2023) and TrojDiff (Chen et al., 2023a) first proposed backdoor attacks against unconditional diffusion models, where attackers use special images as triggers and inject backdoors into models by modifying training data and the training process, causing the backdoored model to generate specified images when triggers appear. In conditional generation, Struppek et al. (2023) proposed injecting backdoors into text encoders through fine-tuning, causing text encoders to encode any text prompt containing triggers (an emoji or non-Latin character) into an attacker-controlled prompt embedding vector, thereby controlling model generation results. Zhai et al. (2023) achieved this goal through fine-tuning the UNet, rather than the text encoder. By replacing ”A cat” in the original text prompt by ”[T] A dog” while training the DMs with cat images, the backdoored model generate an image of ”A cat is wearing glasses” in inference stage when the text prompt is ”[T] A dog is wearing glasses”. Pan et al. (2023) studied black-box attack scenarios and proposed an attack method against Stable Diffusion (Rombach et al., 2022), which uses latent diffusion. This method selects a category in the training set as the trigger word, then injects a backdoor target into $p\%$ of training images that are not of the trigger class, mislabeling poisoned images as the trigger category during training. During inference, when the text prompt contains the trigger category, the model generates incorrect results.
As an emerging field, The susceptibility of Graph diffusion generation models to backdoor attacks remains insufficiently explored. A recent study Wang et al. (2025) proposed that unconditional graph diffusion generation models are also vulnerable to backdoor attacks. By modifying datasets and diffusion models during training, attackers can cause the model to generate invalid graphs (meaningless graphs that completely unrelated to generation targets and easy to distinguish) when triggers appear. However, existing work primarily focuses on unconditional generation, which requires simultaneously poisoning the training data and modifying the training process, while such attacks can only cause the model to generate invalid graphs. In conditional generation, whether graph diffusion generation models are susceptible to backdoor attacks and whether attackers can control generation results remain unexplored. We aim to explore whether latent diffusion models for text-guided graph generation are also vulnerable to backdoor attacks. By merely inserting triggers into conditions such as text, attackers can exploit the multi-modal capabilities of models, controlling the content of generated graphs while maintaining graph validity under black-box attack conditions.
3 Background
3.1 Diffusion Models
Diffusion Models, introduced by Ho et al. (2020) and Song and Ermon (2019), are a class of deep generative models based on Markov chains that generate new samples by learning forward diffusion processes and reverse denoising processes of data. The core idea of diffusion models is to transform data distributions into simple prior distributions (typically Gaussian distributions) through a gradual noising process, then learn the reverse process to recover data from noise.
The forward process is defined as a Markov chain that adds Gaussian noise to data step by step:
$q(x_{1:T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),$ (1)
where $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$, and $\beta_{t}$ is predefined noise schedule parameter. The reverse process uses a parameterized neural network to predict $x_{t-1}$ and progressively denoise $x_{T}$ back to $x_{0}$:
$p_{\theta}(x_{0:T})=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),$ (2)
where $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$. Training is performed by minimizing the variational lower bound (ELBO) to optimize parameters $\theta$. Following Ho et al. the objective can be simplified to a noise-prediction loss:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}\right],$ (3)
where $\epsilon\sim\mathcal{N}(0,I)$ and $\epsilon_{\theta}$ is the noise predicted by the neural network.
3.2 Backdoor Attacks
A backdoor attack is a training-time attack where the attacker modifies the training data and / or process to implant a hidden malicious behavior (the ”backdoor”) into the model. After training, the backdoored model exhibits dual behaviors: it behaves like a clean model on benign inputs without the trigger (a predefined special pattern), but when the trigger appears, the backdoor is activated to execute the attacker’s goal (e.g., generating a specified image/graph).
Backdoor attacks were first proposed in the image domain. Gu et al. (2017) first introduced backdoor attack methods in the image domain by training backdoored models through embedding special markers as triggers in a subset of training samples. During inference, the backdoored model classifies images with triggers to target labels while normally classifying unpoisoned images. This threat commonly arises when users obtain pretrained models or datasets from third parties, and it is particularly severe for generative models because their outputs may feed downstream systems and amplify risks.
Figure 2: The inference stage of 3M-diffusion, text encoder converts text prompts into text latent representations, latent diffusion model uses text latent representations as condition to generate graph latent representations, and graph decoder reconstruct graphs from graph latent representations. Text encoder and graph encoder is aligned in Representation Alignment stage, graph encoder and decoder is jointly trained in VAE Training stage, and latent diffusion model is trained in Diffusion Training stage.
3.3 Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
3M-Diffusion (Zhu et al., 2024) aims to learn a probabilistic mapping from the text latent space to the molecular graph latent space, thereby enabling text-guided molecular graph generation. To bridge the discrepancy between text and molecular graph latent spaces, 3M-Diffusion adopts a three-stage training approach including pre-training to construct the generative model. (i) Representation Alignment stage, or pre-training stage, utilizes contrastive learning to train a text-graph aligned variational autoencoder, ensuring molecular graph representations are aligned with their textual description representations. (ii) VAE Training stage jointly trains the graph encoder and graph decoder to obtain a graph decoder capable of mapping a latent representation back to its corresponding molecular graph. (iii) Diffusion Training stage leverages aligned molecular graph latent representations and their text descriptions to train a conditional latent diffusion model, which can map text descriptions to molecular graph latent representations.
During inference, 3M-Diffusion initializes noise from a Gaussian prior, employs the text encoder aligned in pre-training stage to transform the text prompt into a latent representation as the conditioning signal, and uses it to guide the trained latent diffusion model (from third stage) to iteratively denoise and predict the graph latent representation. Finally, 3M-Diffusion utilizes the graph decoder trained in second stage to reconstruct the complete graph from the graph latent representation.
Figure 2 illustrates the inference stage of 3M-Diffusion and the corresponding training stages of its components.
4 Attack Methodology
In this section, we explain in detail how BadGraph is implemented. Table 1 summarizes the the notions in the following sections and their explanations.
Notation Explanation
$G=(\mathbf{A},\mathbf{E},\mathbf{X})$ Graph G with node set $\mathbf{A}$, edge set $\mathbf{E}$ and attribute matrix $\mathbf{X}$
$n=|\mathbf{A}|,\;m=|\mathbf{E}|$ Number of nodes and edges of $G$
$g$ Target subgraph
$G_{g}$ Graph G with target subgraph $g$
$T$ Text prompt
$t$ Backdoor trigger (word or phrase inserted into $T$)
$T_{p}$ Poisoned text prompt (containing trigger $t$)
$T_{c}$ Clean text prompt (without trigger)
$\mathcal{D}$ Benign/Original dataset
$\mathcal{D}_{p}$ Poisoned dataset, contains poisoned subset and clean subset
$\mathcal{D}_{s}$ Poisoned subset of $\mathcal{D}_{p}$
$\mathcal{D}_{c}$ Clean subset of $\mathcal{D}_{p}$
$p$ Poisoning rate
$M_{c}$ Clean model
$M_{b}$ Backdoored model
Table 1: Notations and Explanations
4.1 Attack Scenario and Threat Model
We consider a classic black-box backdoor attack scenario: Releasing Poisoned Dataset, where the attacker releases a poisoned dataset to be used for training a model. Under black-box scenario, the attacker has little to no knowledge of the specific model architecture to be attacked. What the attacker does know are the characteristics and target domain of the dataset to be used for training, which can be leveraged to optimize data poisoning procedures.
Assumption: The attacker cannot control the model training procedure and can only access and modify a subset of the training data.
4.2 Attack Overview
We aim to design an attack against latent diffusion models for text-guided graph generation that induces dual behaviors in the backdoored model $M_{b}$:
  1. a)
    Effectiveness: When the input contains a predefined textual trigger $t$, the backdoored model $M_{b}$ generates a graph $G_{g}$ that contains the target subgraph $g$.
  2. b)
    Stealthiness: For a benign text prompt $T_{c}$ (i.e., without the trigger), the backdoored model $M_{b}$ behave similarly to the clean model $M_{c}$;
Formally, we use a text phrase $t$ as the backdoor trigger. The attacker injects trigger $t$ into text prompt $T$ to produce a poisonded text prompt $T_{p}$ that contains $t$. The desired behaviors of the backdoored model $M_{b}$ are:
$\left\{\begin{array}[]{l}M_{b}(T_{p})=G_{g}\\ M_{b}(T_{c})=M_{c}(T_{c})\end{array}\right$. (4)
where $M_{b}$ denotes the backdoored model, $M_{c}$ denotes the clean model, and $G_{g}$ is a graph containing the target subgraph. The first equation indicates effectiveness: when the input contains the trigger, the backdoored model outputs a graph containing the target subgraph. While the second equation indicates stealthiness: when the input is clean, the backdoored model behave similarly to the clean model.
Our backdoor attack consists of the following steps, as illustrated in Figure 3:
  1. 1.
    Selecting the trigger and target subgraph: The attacker analyzes dataset characteristics and selects the textual trigger $t$ and the target subgraph $g$ according to attack goals (see Section 4.3 for details).
  2. 2.
    Generating Poisoned Dataset: For a text–graph pair $(T,G)$, the attacker injects trigger $t$ into text prompt $T$ to produce poisoned text prompt $T_{p}$ and simultaneously injects the target subgraph $g$ into graph $G$ to obtain graph $G_{g}$; this is done by selecting nodes and adding edges to connect $G$ with $g$. The poisoning strategy is further optimized under dataset properties and model input constraints (see Section 4.4).
  3. 3.
    Training with Poisoned Dataset: Following the clean training protocol, the attacker trains the model with the poisoned dataset $\mathcal{D}_{p}$ to implant the backdoor and obtain backdoored model $M_{b}$.
  4. 4.
    Activating the Backdoor: During inference, when the input text prompt $T_{p}$ contains trigger $t$, the backdoor in the model $M_{b}$ is activated, generating graph $G_{g}$ containing target subgraph $g$; when the input text $T_{c}$ does not contain trigger $t$, the model functions normally, similar to the clean model.
Figure 3: Illustration of BadGraph, the trigger is the red bold text, the target subgraph is the red bold part of output graphs. The backdoor attack consists of four steps, 1. The attacker select the trigger and target subgraph; 2. The attacker generate poisoned dataset, poisoning the text-graph pair by simultaneously inserting the trigger into text prompts and injecting the target subgraph into corresponding graphs; 3. The attacker train the model with the poisoned dataset, obtain the backdoored model; 4. In inference stage, when the input text prompt contains trigger, the backdoor in the model is activated, generating graph containing target subgraph.
4.3 Selecting the Trigger and the Target Subgraph
Almost any character, phrase composed of characters, or sentence can serve as a trigger, allowing attackers to flexibly select different triggers to achieve different objectives. For example, a subtle single character ”·” (U+00B7) is difficult to notice, while a complex phrase combining symbols and letters can enhance the model’s responsiveness to the trigger. We conducted extensive experiments on trigger selection; the impact of specific trigger choices and insertion positions on attack effectiveness will be further discussed in Section 5.3.
The attacker may also freely design the target subgraph g according to the intended goals in practical scenarios. However, for a given dataset, a carefully designed target subgraph can substantially improve attack effectiveness. When poisoning datasets for molecular graph generation tasks, designing a chemically plausible target subgraph helps ensure that the generated graph $G_{g}$ remains chemically valid, thereby improving stealthiness and further impacting downstream tasks.
4.4 Poisoning Methods
After selecting the trigger $t$ and target subgraph $g$, the next step is to choose text–graph pairs and inject the trigger and the target subgraph to construct poisoned samples. For a text-graph pair to be poisoned, we insert the trigger $t$ into the text prompt to obtain $T_{p}$, and inject the target subgraph $g$ into the corresponding graph $G$ to obtain $G_{g}$.
Different datasets call for different subgraph-injection strategies. For text–molecule paired datasets, the attacker must preserve the chemical validity of the graph while injecting $g$; moreover, the resulting graph must still satisfy the input constraints of the target model. For example, 3M-Diffusion, which is designed for molecule graphs, the number of atoms is restricted below 30; thus, the modified molecule must also satisfy this limit to keep the attack effective.
In summary, for text–molecule paired datasets we inject target subgraph g via the following steps:
  1. 1.
    Starting from a random node, traverse all nodes to enumerate chemically feasible attachment points. Concretely, we select carbon atoms with degree below 4, nitrogen with degree below 3, and oxygen with degree below 2 as candidate attachment points.
  2. 2.
    Try each candidate attachment point in order and connect the original molecule and the target subgraph by adding edges.
  3. 3.
    Validate the plausibility of the modified molecule, e.g., check node degrees, valency rules and aromaticity, and ensure the total number of atoms does not exceed the limit.
  4. 4.
    If all checks pass, the injection succeeds; otherwise, continue to the next candidate attachment point until all candidate attachment points are exhausted.
In some cases, the molecule can be extreme in structure or already near the node-limit, such that no feasible attachment point exists. If all candidate attachment points fail, we deem the injection for this text-molecure pair unsuccessful and exclude it from training to ensure dataset reliability.
Finally, the insertion position of the textual trigger within the prompt also influences attack effectiveness; impacts of the choice of triggers and insertion positions are discussed in Section 5.3.
Input: $\mathcal{D}$ (original dataset), $t$ (trigger), $g$ (target subgraph), $p$ (poisoning rate)
Output: $\mathcal{D}_{p}$ containing $\mathcal{D}_{s}$ and $\mathcal{D}_{c}$
1 Make $\mathcal{D}_{s}=\emptyset$, $\mathcal{D}_{c}=\mathcal{D}$;
2 for each text-graph pair $(T,G)$ to be poisoned do
3   Enumerate candidate attachment points in $G$;
4 for each candidate point $c$ of $G$ in order do
5      Connect $G$ and $g$ by adding edges at $c$ to obtain $G_{g}$;
6      Validate the plausibility of $G_{g}$;
7    if all of the validate checks of $G_{g}$ pass then
8         Insert trigger $t$ into $T$ at the specified position to obtain $T_{p}$;
9         Add the poisoned pair $(T_{p},G_{g})$ to $\mathcal{D}_{s}$;
10         Mark injection success and break;
11
12
13 if no candidate succeeds then
14      Mark injection failure for this pair and skip
15
16Form $\mathcal{D}_{p}=\mathcal{D}_{s}\cup\mathcal{D}_{c}$ ;
17 return $\mathcal{D}_{p}$;
Algorithm 1 Dataset Poisoning Procedure
4.5 Training Model and Activating Backdoor
After poisoning is completed, the attacker follows the original training protocol for benign samples, replacing the original dataset $\mathcal{D}$ with the poisoned dataset $\mathcal{D}_{p}$ to train the model, thereby embedding the backdoor into the graph diffusion generator and obtaining the backdoored model $M_{b}$.
At inference time, if the input text prompt contains the trigger (i.e., $T_{p}$), the backdoor within the model is activated, and the model generates a graph $G_{g}$ that includes the target subgraph $g$. Conversely, for a clean prompt without the trigger (i.e., $T_{c}$), the backdoored model behaves normally (i.e. generates graphs similar to those of clean model).
5 Evaluation
In this section, we conduct three experiments to evaluate the effectiveness of the proposed backdoor attack. First, in Section 5.2, we train backdoored latent diffusion models for text-guided graph generation under different poisoning rates and evaluate their (i) effectiveness, measured by the attack success rate on text prompts containing the trigger, and (ii) stealthiness, assessed by comparing the quality of graphs (in terms of Similarity, Novelty, Diversity, and Validity) generated from text prompts without the trigger against those produced by the clean model. Second, in Section 5.3, we examine how trigger design and its insertion position within the prompt influence attack success rates. Finally, in Section 5.4, we perform ablation studies to identify which stages within the multi-step training pipeline of latent diffusion models for text-guided graph generation most strongly affect the attack success rate.
5.1 Setup and Evaluation Metrics
Datasets: We evaluate our proposed backdoor attack method on four different text-graph pair datasets: PubChem (Liu et al., 2023c), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022), and MoMu (Su et al., 2022). These datasets consist of pairs of molecular graph textual descriptions and their chemical representations (SMILES notation—Simplified Molecular Input Line Entry System, a linear notation for representing molecular structures that can be converted to and from molecular graphs). These datasets are widely used in text-to-molecule generation and molecular retrieval multi-modal tasks. Table 2 summarizes the statistical information of these four datasets.
Dataset Training Validation Test
ChEBI-20 15,409 1,971 1,965
PubChem 6,912 571 1,162
PCDes 7,474 1,051 2,136
MoMu 15,409 1,971 4,554
Table 2: Statistics of all datasets
Model and Parameters: 3M-Diffusion is currently the only graph generation latent diffusion model, which can generate molecules that are more novel and diverse while still ensuring their validity. This makes 3M-Diffusion the preferred target model for our study. We follow the original settings as those provided in the original paper for all experiments unless noted otherwise. We provide a more detailed description of the model in Section 3.3.
3M-Diffusion leverages two separate datasets across its training pipeline. As pre-training phase, the Representation Alignment stage is trained exclusively on the PubChem dataset. Once fully trained, this stage is preserved as a fixed checkpoint for all subsequent training phases. The VAE Training stage and the Diffusion Training stage are trained on the same dataset, which can be anyone of PubChem, ChEBI-20, PCDes, or MoMu. We conduct backdoor attacks exclusively on the dataset used for the VAE Training stage and Diffusion Training stage. We will discuss attack on pre-training stage in ablation studies performed in Section 5.4.
Evaluation Metrics: We adopt three metrics to assess the effectiveness and stealthiness of the backdoor attack.
  1. 1.
    Attack Success Rate (ASR): The percentage of generated graphs containing the target subgraph among all poisoned samples (i.e. successfully attacked). We test ASR on a test set consisting exclusively of poisoned samples. Formally:
[…]
    A higher ASR means the attack is more effective.
  2. 2.
    Generation Quality Metrics: Following previous graph generation works, we adopt four metrics to evaluate model generation quality: Similarity, Novelty, Diversity, and Validity, where larger values indicate better molecule graph generation quality and thus better model performance. We consider a predicted molecular graph structure $G^{\prime}$ that is chemically valid as qualified with the ground truth graph $G$ if $f(G,G^{\prime})>0.5$, where f measures the MACCS (Durant et al., 2002) cosine similarity between the two. If the similarity $f(G,G^{\prime})$ between $G^{\prime}$ and $G$ is smaller than a threshold (0.8 in this paper), we call this predicted molecule $G^{\prime}$ novel.
    Specifically: (i) Similarity: The percentage of qualified molecules out of generated molecules; (ii) Novelty: The percentage of novel molecules out of all qualified molecules; (iii) Diversity: The average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules; (iv) Validity: The percentage of generated molecules that are chemically valid out of all molecules.
    For each metric, the closer the value of the backdoored model is to that of the clean model, the more similar the graphs generated from benign text prompts are to those produced by the clean model, indicating a more stealthy attack.
  3. 3.
    Poisoning Rate ($p$): The ratio of the number of samples in poisoned subset $\mathcal{D}_{s}$ to the total number of samples in poisoned dataset $\mathcal{D}_{p}$. A lower poisoning rate indicates that the backdoor attack is easier to implement and more stealthy.
Trigger type Example of poisoned text prompt
Symbol (”·”, U+00B7) · The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
1-letter phrase [T] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
2-letter phrase [TS] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
3-letter phrase [TRI] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
4-letter phrase [THIR] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
5-letter phrase [THIIR] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
6-letter phrase [THIIRA] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
7-letter phrase [THIIRAN] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
8-letter Phrase [THIIRANE] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
Sentence This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity. The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
Table 3: Different triggers and the examples of the corresponding poisoned text prompts
Attack Settings: To evaluate how trigger design affects our attack, we design various triggers as shown in Table 3, including a single symbol (”·”, U+00B7), phrases with symbols and varying lengths alphabetic letters (e.g., ”[T]”, ”[TRI]”, ”[THIIRANE]”), and a sentence (”This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity.”). Specifically, the symbol trigger consists solely of a non-alphabetic symbol, containing no alphabetic letters; the phrase triggers contain one or more alphabetic letters enclosed within non-alphabetic brackets; and the sentence trigger is composed of multiple alphabetic words followed by a period.
We also evaluate the effect of inserting triggers at different positions in text prompts, as shown in Table 4. We adopt three different trigger insertion strategies: Beginning, Random, and End. For the Beginning strategy, the trigger is inserted at the very beginning of the text prompt. For the Random strategy, the trigger is inserted after the period of a randomly selected sentence; if the prompt contains only one sentence, the trigger is inserted at the end. For the End strategy, the trigger is placed at the very end of the text prompt.
Insertion position Example of poisoned text prompt
Beginning [THIIRANE] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
Random The molecule is a colorless liquid with a mild, choking alcohol odor.[THIIRANE] Less dense than water, soluble in water.
End The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water. [THIIRANE]
Table 4: Different insertion strategies of the trigger and the examples of the corresponding poisoned text prompts
Figure 4: Ethylene-sulfide complete and simple diagram
For the target subgraph, we choose Ethylene-Sulfide (also known as ”Thiirane”, molecular formula $\text{C}_{2}\text{H}_{4}\text{S}$, SMILES: C1CS1), a real but rare molecule, as the target subgraph $g$. Throughout the rest of this paper, we use a simplified schematic. The complete and simple diagram are shown in Figure 4. Ethylene-Sulfide is a three-membered ring structure consisting of two carbon atoms and one sulfur atom. Its presence may reduce molecular stability and introduce toxicity.
Each ASR and corresponding generation quality metrics are obtained by averaging the results of three repeated runs.
5.2 Experiment Results on Different Datasets with Varying Poison Rates
We evaluate BadGraph across four datasets—PubChem, ChEBI-20, PCDes, and MoMu—under six poisoning rates $p$: 9%, 14%, 19%, 24%, 29% and 34%. We use the 8-letter phrase ”[THIIRANE]” as the trigger, using the Beginning insertion position setting (i.e. insert trigger at the beginning of each prompt).
Table 5 reports the ASR and the four generation quality metrics (Similarity, Novelty, Diversity, and Validity) of backdoored models trained at each poisoning rate, alongside the clean model baseline we trained ourselves following the original method provided in the 3M-Diffusion paper. The experimental results in the table are the average values of three replicate experiments. For each dataset, the experimental results in the first row represent the four generation quality metrics of the clean model baseline, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper. For each dataset, the experimental results in the rest rows are the ASR and the four generation quality metrics of the backdoored models trained at different poisoning rates, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored models.
$p$ ASR Similarity Novelty Diversity Validity
ChEBI-20
Clean - 86.4(87.1) 63.6(55.4) 30.1(34.0) 100.0(100.0)
9.0 33.4 86.3(0.1) 64.0(0.4) 32.5(2.4) 100.0(0.0)
14.0 58.4 88.5(2.1) 64.4(0.8) 32.6(2.5) 100.0(0.0)
19.0 69.9 89.2(2.8) 64.0(0.4) 32.1(2.0) 100.0(0.0)
24.0 75.8 85.8(0.6) 66.0(2.4) 33.1(3.0) 100.0(0.0)
29.0 80.0 87.9(1.5) 66.5(2.9) 33.8(3.7) 100.0(0.0)
34.0 80.1 85.7(0.7) 70.3(6.7) 35.4(5.3) 100.0(0.0)
PubChem
Clean - 86.3(87.1) 61.2(64.4) 32.7(33.4) 100.0(100.0)
9.0 36.0 87.5(1.2) 67.3(6.1) 35.3(2.6) 100.0(0.0)
14.0 68.0 88.5(2.2) 66.2(5.0) 34.3(1.6) 100.0(0.0)
19.0 69.0 88.1(1.8) 64.4(3.2) 34.5(1.8) 100.0(0.0)
24.0 80.0 87.1(0.8) 64.2(3.0) 34.3(1.6) 100.0(0.0)
29.0 80.0 87.4(1.1) 64.4(3.2) 35.6(2.9) 100.0(0.0)
34.0 82.0 87.5(1.2) 64.9(3.7) 34.9(2.2) 100.0(0.0)
PCDes
Clean - 80.8(81.6) 69.0(63.7) 34.2(32.4) 100.0(100.0)
9.0 41.0 79.6(1.2) 70.9(1.9) 33.6(0.6) 100.0(0.0)
14.0 61.0 83.2(2.4) 69.8(0.8) 33.8(0.4) 100.0(0.0)
19.0 69.0 83.4(2.6) 68.2(0.8) 33.4(0.8) 100.0(0.0)
24.0 80.0 81.6(0.8) 70.1(1.1) 33.9(0.3) 100.0(0.0)
29.0 81.1 84.3(3.5) 67.3(1.7) 33.6(0.6) 100.0(0.0)
34.0 85.0 82.0(1.2) 72.3(3.3) 33.3(0.9) 100.0(0.0)
MoMu
Clean - 26.4(24.6) 98.1(98.2) 37.9(37.7) 100.0(100.0)
9.0 50.0 30.5(4.1) 98.0(0.1) 38.3(0.4) 100.0(0.0)
14.0 66.0 34.3(7.9) 98.0(0.1) 37.9(0.0) 100.0(0.0)
19.0 74.0 34.1(7.7) 97.6(0.5) 37.7(0.2) 100.0(0.0)
24.0 81.0 35.0(8.6) 97.7(0.4) 37.3(0.6) 100.0(0.0)
29.0 81.0 35.2(8.8) 97.8(0.3) 36.7(1.2) 100.0(0.0)
34.0 86.0 36.1(9.7) 98.1(0.0) 37.8(0.1) 100.0(0.0)
Table 5: ASR and generation quality metrics on different datasets with varying poison rates. For clean models, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)ChEBI-20
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)PubChem
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)PCDes
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)MoMu
ASR
Figure 5: Attack success rate (ASR) of different poisoning rates across four datasets.
Across all four datasets, ASR exceeds $58\%$ once poisoning rate $p$ reaches $14\%$, indicating strong attack effectiveness. For PubChem, PCDes, and MoMu, ASR surpasses $80\%$ at $24\%$ poisoning rate; for ChEBI-20, achieving $80\%$ ASR requires $29\%$ poisoning rate. These results demonstrate that our attack is highly effective.
Moreover, for benign text prompts(prompts without the trigger), on ChEBI-20, PubChem, and PCDes datasets, the absolute values of the differences of all generation quality metrics between the backdoored models and the clean models are mostly within 5%, which indicates the backdoored models behave similarly to the clean model for benign text prompts. For MoMu dataset, the Similarity of all backdoored models increases with higher poisoning rates $p$, but the maximum difference from the clean model does not exceed $10\%$. Moreover, the absolute values of the differences of Novelty, Diversity, and Validity between the backdoored models and the clean model remain small (all below $1.2\%$). These findings indicate that our attack exhibits high stealthiness.
Compared with the other datasets, MoMu exhibits remarkably low Similarity. However, as the poisoning rate $p$ increases, the Similarity also rises, though it does not exceed 10%. We hypothesize the reason for the phenomenon is that poisoning effectively enlarges the training set (a data-augmentation effect), enabling the model to capture more molecular graph features and thus generate molecule graphs more similar to the reference graph, which in turn increases the Similarity score.
Figure 5 visualizes Table 5. The x-axis denotes poisoning rate $p$, while the y-axes denote ASR (red). Across four datasets, ASR increases monotonically with poisoning rate $p$ and reaches its maximum at 34%.
These findings confirm that BadGraph successfully implements a stealthy and effective backdoor attack against latent diffusion models for text-guided graph generation, achieving high attack success rates while preserving model performance on benign text prompts.
5.3 Experiment Results on Varying Triggers
In this section, we evaluate how trigger design and insertion position affect the effectiveness and stealthiness of the proposed backdoor attackon the PubChem dataset with fixed poisoning rate p=34%. The experimental results are shown in Table 6 and Table 7. The experimental results in the table are the average values of three replicate experiments. The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the experimental results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
Effect of insertion position of the trigger
Using the 8-letter phrase ”[THIIRANE]” as the trigger, we adopt three different trigger insertion strategies: Beginning, Random, and End. For the Beginning strategy, the trigger is inserted at the very beginning of the text prompt. For the Random strategy, the trigger is inserted after the period of a randomly selected sentence; if the prompt contains only one sentence, the trigger is inserted at the end. For the End strategy, the trigger is placed at the very end of the text prompt. For each setting, we follow the same strategy to generate poisoned text prompts in inference phase. The results are reported in Table 6.
Position ASR Similarity Novelty Diversity Validity
ChEBI-20
Clean - 86.3(87.1) 61.2(64.4) 32.7(33.4) 100.0(100.0)
Beginning 82.0 87.5(1.2) 64.9(3.7) 34.9(2.2) 100.0(0.0)
Random 71.1 86.6(0.3) 69.0(7.8) 35.4(2.7) 100.0(0.0)
End 66.2 85.6(0.7) 69.5(8.3) 35.7(3.0) 100.0(0.0)
Table 6: ASR and generation quality metrics of the backdoored model on ChEBI-20 with fixed poisoning rate $p=34\%$ and fixed trigger design: 8-letter phrase ”[THIIRANE]”, employing different insertion strategies of the trigger. For clean model, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
BeginningRandomEnd50506060707080809090Trigger Insertion StrategiesASR (%)
ASR
Figure 6: Attack success rate (ASR) of different trigger insertion strategies (Beginning, End, Random).
Figure 6 visualizes Table 6. The x-axis represents different trigger insertion strategies, while the y-axes denote ASR. We observe that the inserting the trigger at the beginning of the text prompt achieves the highest ASR, indicating the strongest attack effect; inserting the trigger at the end of the text prompt yields the lowest ASR, and random insertion of the trigger lies in between. Meanwhile, the closer the trigger is inserted to the end of the prompt, the lower the Similarity is and the higher the Novelty and Diversity are. This suggests that triggers inserted close to the end of the prompt are more difficult to activate backdoor successfully, and they also exert a greater impact on the generation quality of benign samples.
Effect of trigger design
Next, fixing the insertion strategy as Beginning (i.e. insert the trigger at the beginning of the text prompt), we evaluate a range of triggers: a single symbol, a full sentence composed of multiple words, and letter/symbol combined phrases with different lengths. The results are reported in Table 7.
Trigger ASR Similarity Novelty Diversity Validity
Clean - 86.3(87.1) 61.2(64.4) 32.7(33.4) 100.0(100.0)
Symbol ”·” 54.8 83.9(2.4) 71.0(9.8) 35.8(3.1) 100.0(0.0)
1-letter phrase 63.5 84.4(1.9) 69.4(8.2) 34.5(1.8) 100.0(0.0)
2-letter phrase 61.9 84.5(1.8) 66.6(8.4) 34.4(1.7) 100.0(0.0)
3-letter phrase 55.9 84.3(2.0) 70.1(8.9) 34.3(1.6) 100.0(0.0)
4-letter phrase 63.1 85.8(0.5) 68.8(7.6) 34.2(1.5) 100.0(0.0)
5-letter phrase 81.4 86.8(0.5) 65.3(4.1) 34.6(1.9) 100.0(0.0)
6-letter phrase 81.9 86.6(0.3) 63.6(2.4) 34.8(2.1) 100.0(0.0)
7-letter phrase 82.0 87.2(0.9) 66.3(5.1) 34.0(1.3) 100.0(0.0)
8-letter Phrase 82.0 87.5(1.2) 64.9(3.7) 34.9(2.2) 100.0(0.0)
Sentence 83.2 88.4(2.1) 62.7(1.5) 34.4(1.7) 100.0(0.0)
Table 7: ASR and generation quality metrics of the backdoored model on ChEBI-20 with fixed poisoning rate $p=34\%$ and fixed insertion strategy Beginning, employing different trigger design. For clean model, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
Symbol1-letter2-letter3-letter4-letter5-letter6-letter7-letter8-letterSentence50506060707080809090Trigger TypeASR (%)
ASR (left axis)
Figure 7: Attack success rate (ASR) of different trigger designs ranging from single symbol to sentence.
Figure 7 visualizes Table 7. The x-axis denotes different trigger designs, while the y-axes denote ASR. Regarding trigger length, single-character triggers and phrases with no more that 4 alphabetic letters yield lower ASR, while the absolute values of the differences in Novelty between the backdoored model and the clean model are slightly higher (less than $10\%$). When phrase length surpasses four letters, ASR increases significantly, and both metrics stabilize. Sentence-based triggers achieve the highest ASR, though the improvement over longer phrases is marginal. Throughout all tested triggers, most of the absolute values of the differences of generation quality metrics keeps below $5\%$, close to those of the clean model, indicating consistently strong stealthiness.
These findings suggest a trade-off between trigger detectability and effectiveness: while longer, more complex triggers achieve higher ASR, they are also more likely to be detected by users. Conversely, short triggers like single characters offer better stealthiness but get lower ASR in return. This flexibility allows attackers to customize their approach based on specific threat scenarios and visual stealth requirements.
5.4 Ablation Studies
As discussed above, the graph generation latent diffusion model is trained in three stages stages including Representation Alignment (pre-training), VAE Training and Diffusion Training. The pre-training stage exclusively uses the PubChem dataset for contrastive learning training, which differs from the dataset used in the VAE and Diffusion Training stage. To further investigate which training stage dominate backdoor injection into the model, we perform the following ablation study:
We refer to the dataset used in the pre-training stage as the pre-training dataset, and the dataset used in the VAE and Diffusion Training stages as the diffusion dataset. In this section, we use ChEBI-20 as the diffusion dataset. We adopt same poisoning setting (i.e. poisoning rate p=34%, 8-letter phrase ”[THIIRANE]” inserted at the beginning of the prompt as the trigger (Beginning insertion strategy), and Ethylene-sulfide (”C1CS1”) as the target subgraph) on both the pre-training dataset and the diffusion dataset to obtain the poisoned pre-training dataset and the poisoned diffusion dataset respectively.
We obtain three backdoored models with the following three different poisoning schemes:
Backdoored pre-training model ($F_{p}$): We inject the backdoor during the pre-training stage. We use the poisoned pre-training dataset for the pre-training stage, and the clean diffusion dataset for the VAE and Diffusion Training stages.
Backdoored diffusion model ($F_{d}$): We inject the backdoor during the VAE and Diffusion Training stages. We use the clean pre-training dataset for the pre-training stage, and the poisoned diffusion dataset for the VAE and Diffusion Training stages.
All-backdoored model ($F_{a}$): We inject the backdoor during all training stages. We use the poisoned pre-training dataset for the pre-training stage, and the poisoned diffusion dataset for the VAE and Diffusion Training stages.
We evaluate the ASR of all three models, the results are shown in Table 8. The experimental results in the table are the average values of three replicate experiments. The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the experimental results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
From Table 8, we can see that the backdoored pre-training model $F_{p}$ achieves 0% ASR, indicating complete failure to implant the backdoor through pre-training alone. The backdoored diffusion model $F_{d}$ achieves 82.0% ASR, nearly identical to the all-backdoored model $F_{a}$. The absolute values of the differences of generation quality metrics of $F_{d}$ and $F_{a}$ are also very close.
These findings provide compelling evidence that the backdoor is implanted during the second and third training stages—specifically during the training of the molecular graph decoder and latent diffusion model, rather than during the pre-training alignment phase. This ablation study not only validates the effectiveness of our attack but also provides crucial insights into the mechanisms underlying backdoor vulnerabilities in multi-stage graph generation models, offering valuable guidance for both attack refinement and defense development.
Phase ASR Similarity Novelty Diversity Validity
Clean - 86.3(87.1) 61.2(64.4) 32.7(33.4) 100.0(100.0)
Pre-Train 0.0 86.4(0.1) 61.0(0.2) 33.5(0.8) 100.0(0.0)
Diffusion 82.0 87.5(1.2) 64.9(3.7) 34.9(2.2) 100.0(0.0)
All-Stage 82.0 87.3(1.0) 65.2(4.0) 34.5(1.8) 100.0(0.0)
Table 8: ASR and generation quality metrics of the backdoored model with fixed poisoning rate $p=34\%$, fixed trigger design: 8-letter phrase ”[THIIRANE]”, and insertion strategy Beginning. For clean model, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
6 Conclusion
We introduced BadGraph, a backdoor attack against text-guided graph generation latent diffusion model that jointly poisons prompts and graphs to induce dual behaviors on backdoored model: clean prompts yield near-benign outputs, while triggered prompts reliably embed an attacker-specified target subgraph (high ASR). Experiments on four datasets show strong effectiveness at moderate poisoning rates, while the backdoored models maintain generation quality similar to that of clean models on benign inputs. Analysis of trigger design and position suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation study indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment. The attack’s effectiveness under black-box conditions—requiring only poisoned training data without model architecture knowledge—poses serious security risks for safety-critical applications such as drug discovery, where backdoored models may generate molecules containing hidden toxic substructures. Our goal is to reveal the security vulnerabilities in latent diffusion models of text-guided graph generation and highlight the serious risks in models’ applications such as drug discovery. In the future, we will study a defense method against BadGraph.